INFRARED PHOTODISSOCIATION SPECTROSCOPY OF METAL ION WATER COMPLEXES

\underline{B. BANDYOPADHYAY}, P. D. CARNEGIE and M. A. DUNCAN, University of Georgia, Athens, Georgia-30605, USA.

Metal ion-water complexes are produced in a supersonic expansion cluster source via laser vaporization technique. Infrared photodissociation spectroscopy has been performed in the O-H stretch region. DFT calculations have also been carried out to obtain the structures and vibrational frequencies. Infrared spectra show partially resolved rotational structures which will be analyzed.